OpenStudy (geekfromthefutur):

I will give medal and fan math isnt my best subject. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+68=7x−16241(3x+68)=241(7x−16) 3(3x+6)=4(7x−1)9x+18=28x−19x−9x+18=28x−9x−118=19x−118+1=19x−1+119=19x1=x

OpenStudy (geekfromthefutur):

3x+68 = 7x−16 241(3x+68) = 241(7x−16) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 =28x−9x−1 18+1=19x−1+1 19 = 19x 1=x

OpenStudy (yanasidlinskiy):

Is this the original equation? \[\frac{ 3x+6 }{ 8 }=\frac{ 7x-1 }{ }\]

OpenStudy (geekfromthefutur):

I tired spacing it out thats the best i can do

OpenStudy (yanasidlinskiy):

Was that the original equation what i posted above? or 3x+68 = 7x-16

OpenStudy (geekfromthefutur):

im getting it right now i just took a pic @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Alright:)

OpenStudy (geekfromthefutur):

Does that help?

OpenStudy (yanasidlinskiy):

Ok. Yea it was the first one that i posted. Ok. Even though it's blurry.. replace "x" with 1, and see if the equation is true (left side equals right side) if it is, x=1 is the correct answer.

OpenStudy (geekfromthefutur):

I got another question ima open another one so i can give u another medal

OpenStudy (yanasidlinskiy):

To find the correct answer, we can multiply both sides by 2 to get  now "cross multiply to get to this step 3(3x+6)=4(7x−1) Max got this far, and then distributed the 3 on the left side, and the 4 on the right side: 9x+18=28x−4 (max made a mistake and got 9x+18 = 28x−1 ) can you finish?

OpenStudy (geekfromthefutur):

So the correct answer is 27x ??

OpenStudy (yanasidlinskiy):

Let me check...i think you made a mistake...lol..

OpenStudy (geekfromthefutur):

Dang it

OpenStudy (geekfromthefutur):

Is it 1=x?

OpenStudy (yanasidlinskiy):

Yes. It's x=1. Look: 9x+18 = 28x-1 -9x -9x 18=19x-1 +1 +1 19=19x ------> Divide by 19 x=1!:) You are correct!!:)

OpenStudy (geekfromthefutur):

^__^

OpenStudy (yanasidlinskiy):

Help on anymore? Before I go offline:( Sorry